### v1

# Do graph neural network states contain graph properties?

Abstract
Graph learning models achieve state-of-the-art performance on many tasks, but this often requires increasingly large model sizes. Accordingly, the complexity of their representations increase. Explainability techniques (XAI) have made remarkable progress in the interpretability of ML models. However, the non-relational nature of Graph Neural Networks (GNNs) make it difficult to reuse already existing XAI methods. While other works have focused on instance-based explanation methods for GNNs, very few have investigated model-based methods and, to our knowledge, none have tried to probe the embedding of the GNNs for well-known structural graph properties. In this paper we present a model agnostic explainability pipeline for Graph Neural Networks (GNNs) employing diagnostic classifiers. This pipeline aims to probe and interpret the learned representations in GNNs across various architectures and datasets, refining our understanding and trust in these models.
1 Introduction
In the last decade, significant progress has been made towards modelling non-Euclidean, graph-structured data (Kipf & Welling, 2017) on the one hand, and on interpreting the predictions of deep neural networks (DNN) on the other hand. We often qualify DNNs as black box as their predictions are not inherently interpretable. Occlusion, gradient, perturbation, layer-wise relevance propagation, and attention mechanisms have been proposed to solve this problem (Zeiler & Fergus, 2013; Denil et al., 2015; Li et al., 2016; Sundararajan et al., 2017). These methods focus on highlighting the importance of different input features. They can, however, not be directly applied on GNNs due to the lack of a regular structure (e.g. vertices can have different degrees). In this case, explaining a prediction means identifying important parts of the relational structure, or input features of nodes. An issue is that finding the explanation is itself a combinatorial problem, making XAI (explainable AI) methods for GNN intractable (Longa et al., 2023a; Ying et al., 2019; Lucic et al., 2022).
Previous surveys (Agarwal et al., 2023; Dai et al., 2022) highlighted the lack of comprehensive, robust and model-agnostic explainability methods. We also identified that there are very few model-level explainability methods. As an alternative to these more traditional XAI methods, we propose to apply probing techniques for graph properties (as developed for Natural Language Processing Giulianelli et al. (2018), Belinkov (2021)) to GNN embeddings. In our pipeline (see fig. 1), we investigate both local properties like betweennes centrality, as well as global properties like average path length. To our knowledge, this is the first work to explore this direction.
Figure 1: An example of the probing pipeline. First, a GNN is trained on a specific task, for example detecting whether a graph contains a grid or house shaped pattern. Then, we extract embeddings from the internal layers of the network. We use these embeddings to train the probing model; in this example a binary classifier which can detect whether the embedding contains predictive information for the diameter of the graph. If a linear probe has good performance ($R^{2}$ score) then there exists a hyperplane in the representation space that separates the inputs based on the property
Findings  [^1] All results and experiments accessible on github:
  • We demonstrate the ability of diagnostic classifiers to effectively highlight known graph-theoretic and domain-specific properties in GNN learned latent representations (fig. 5).
  • We explore how different regularization techniques (none, $L_{2}$ weight decay, dropout) affect the representation of graph properties within the same GNN architecture (fig. 10).
  • We compare how various GNN architectures (GCN, R-GCN, GIN, GAT) differ in their ability to represent graph properties, analyzing whether these differences align with their mathematical frameworks (table 6).
  • We apply this pipeline to a toxicity dataset showing that probed graph properties align with chemical knowledge (table 8) before exploring the pipeline’s inferential power on fMRI datasets, uncovering structural properties that might not yet have been extensively studied (table 23).
2 Background
2.1 Graph Neural Networks
Nowadays we have some theoretical understanding of the representational restrictions and capabilities of Graph Neural Networks (GNNs) with regard to the Weisfeiler-Lehman test (Akhondzadeh et al., 2023). We know that this cannot capture certain graph properties, such as connectivity or triangle-freeness (Franks et al., 2024; Kiefer, 2020; Kriege et al., 2018), due to its reliance on local structure. This constraint is also present in (message passing) GNNs.
Graph Convolutional Network (GCN) (Kipf & Welling, 2017) are GNNs where for a single layer, the node representation is computed as: $\bm{X}^{\prime}=\sigma\left(\tilde{\bm{D}}^{-1/2}\cdot\tilde{\bm{A}}\cdot\tilde{\bm{D}}^{-1/2}\cdot\bm{X}\cdot\bm{W}\right)$. We know that GNNs which rely solely on local information, like the GCN and its relational variant (R-GCN(Schlichtkrull et al., 2018), cannot compute important graph properties, such as girth and diameter or eigenvector centrality Garg et al. (2020). We are therefore also investigating more globally aware networks like GAT (Graph Attention Network) (Veličković et al., 2018) and GIN (Graph Isomorphism Network) (Xu et al., 2019).
GAT makes use of self-attention and is thereby more expressive than the GCN. However, its reliance on feature-dependent weights and structure-free normalisation limits its ability to capture specific structural properties that do not directly depend on edges. This is particularly true for tasks where node features alone are not enough, and global graph structures are crucial (e.g., tasks requiring knowledge of subgraphs or non-local patterns). GIN aggregates node features in a way that mimics the Weisfeiler-Lehman test for graph isomorphism, and with its strong inductive learning capabilities, it is likely to excel at encoding complex graph properties and solving classification tasks.
2.2 Graph properties
Graph theory is a branch of mathematics that studies the properties and relationships of graphs. Graphs can be undirected or directed and analysed through both local and global properties. Local properties like node degree which count the number of connections a node has, identifying highly connected hub nodes, or the clustering coefficient which measures how well a node’s neighbours are interconnected, capturing the local density of connections and giving the node a score, are based on a node with regard to its neighbour. In contrast, global properties such as diameter and characteristic path length assess the overall structure. They indicate how far nodes are from one another and how efficiently information can spread through the network. Global graph properties can be associated with higher level complex systems’ characteristics like the presence of some repeated motifs in the sub structures of the graphs or information-flow properties.
We can distinguish different global properties, basic ones like the number of nodes a graph has, clustering and centrality ones, graph motifs and substructures, spectral and small-world properties.
As an higher-order analysis, the recurrence of specific motifs within network substructures—such as triangles, cliques, or feed-forward loops can be seen as the fundamental building blocks that dictate the system’s functionality and resilience. Small-worldness  [^2] We are using the Small-World Index, $SWI=\left(\frac{L-L_{l}}{L_{r}-L_{l}}\right)\times\left(\frac{C-C_{r}}{C_{l}-C_{r}}\right)$ in our experiment because it provides a more balanced and robust measure of small-world properties. Unlike the Small-World Quotient: $Q=\frac{C/C_{r}}{L/L_{r}}$, which can be sensitive to network size and degree, $SWI$ normalises both the clustering coefficient and average path length with respect to both random and lattice reference graphs. This dual normalisation approach ensures that $SWI$ is less prone to false positives or negatives, making it a more reliable metric for our analysis Neal (2017). , as characterised by Barabási Albert & Barabási (2002), reveal how networks can maintain short path lengths despite their expansive size and sparse connectivity. This kind of higher order properties are very interesting in order to understand how the macroscopic behaviour of complex systems emerges from the intricate interplay of their microscopic components Barabási et al. (2002). For example how diseases spread in social networks, how neurons interact in the brain, or how information propagates through the Internet. GNNs synthesise local topological features into global structures, abstract these representations into higher-order graph attributes. Probing their learnt representations should act as a scalable proxy to investigate how global arrangement and connectivity patterns influence a system’s function. In other terms, by dissecting these learned embeddings, we can possibly delve into the intricate relationships between a network’s macroscopic arrangement and its emergent behaviours.
A Graph $G=(V,E)$, $V$ the set of vertices, $E$ the set of edges, can be analysed through both local and global properties. Local properties (like node degree or clustering coefficient) are based on the neighbors of a node.
In contrast, global properties (such as diameter and characteristic path length) assess the overall graph structure. Global graph properties can be associated with higher level complex systems’ characteristics like the presence of repeated motifs in the graphs or information-flow properties. See the appendix B for a list of local and global properties used in our experiments.
GNNs synthesise local topological features into global structures and then abstract these representations into higher-order graph attributes. Probing their learnt representations should act as a scalable proxy to investigate how global arrangement and connectivity patterns influence a system’s function. Based on the message passing paradigm in GNNs, as layers progress, one would expect an increased abstraction in the selection of graph properties. Initially, local features like node degree dominate, but deeper layers progressively capture more global properties, such as connectivity patterns and centrality.
Through hierarchical pooling or readout mechanisms, GNNs can aggregate node embeddings into a single, global graph-level embedding. Graphs that share structural similarities or patterns of interaction among nodes are organised closely in the embedding space, allowing the model to differentiate between classes of graphs, such as those with and without long paths.
2.3 Probing classifiers
In prior work (Hupkes et al., 2018) probing classifiers have been used for linguistic properties. Here, we adapt them for graph features. Unlike unsupervised techniques such as Principal Component Analysis (PCA) or T-SNE, which are useful to visualise input data with regard to the embedding latent space, we adopt a supervised framework to quantitatively assess how specific properties are encoded within the embedding space of DNNs. Let $g:f_{l}(x)\mapsto\hat{z}$ represent a probing classifier, used to map the learned intermediate representations from the original model $f$ to a specific property $\hat{z}$. The choice of a linear classifier for $g$ is motivated primarily by its simplicity. If a linear probe performs well, it suggests the existence of a hyperplane in the representation space that separates the inputs based on their properties, indicating linear separability.
Another advantage of a simple linear probe is avoiding the risk that a more complex classifier might infer features that are not actually used by the network itself Hupkes et al. (2018). While other non-linear probes have been explored in the literature Belinkov (2021), even studies showing improved performance with complex probes maintain the same logic: $\operatorname{Perf}(g,f_{1},\mathcal{D}_{O},\mathcal{D}_{P})>\operatorname{Perf}(g,f_{2},\mathcal{D}_{O},\mathcal{D}_{P})$ holds across representations $f_{1}(x)$ and $f_{2}(x)$ when evaluated by a consistent probe $g$. This consistency ensures valid comparison, underscoring that if a property can be predicted well by a simple probe, it is likely relevant to the primary classification task.
From an information-theoretic perspective, training the probing classifier $g$ can be viewed as estimating the mutual information between the learned representations $f_{l}(x)$ and the property $z$. This mutual information is denoted as $\mathrm{I}(\mathbf{z};\mathbf{h})$, where $\mathbf{z}$ refers to the property and $\mathbf{h}$ represents the intermediate representations Belinkov (2021).
This supervised approach allows us to define hyperplanes or higher-dimensional decision boundaries that partition the embedding space according to the chosen graph property. The $R^{2}$ score serves as this information-theoretic measure indicating how well the hyperplane divides the inputs in the embedding space. A $R^{2}$ near 1 indicates that the embeddings are highly informative about $\hat{z}$, suggesting that the neural model has internalized this property in a linearly accessible manner.
By defining specific properties that could divide the embedding space and assessing how well the corresponding hyperplanes make the embedding space linearly separable, we gain quantitative insights into the abstract features aggregated within the embeddings. This method moves beyond mere hypothesis generation based on clustering patterns observed through techniques like PCA, providing a rigorous framework for understanding how well the embedding space represents complex graph properties. It can also be thought as complementary from the T-SNE and PCA visualisation techniques, as it provides a quantitative measure of the separability of the embeddings based on hypothesised properties of interest.
The best illustration of this comes with fig. 3. We illustrate the evolution of the separability of graphs in the embeddings also in fig. 5 using a T-SNE visualisation and the corresponding separability with the properties thanks to the probing. This highlight the most interesting results of the paper, showing that the separability of house-only and grid-only graphs in the negative class (purple) match with the presence of the property number of triangles in the 5th layer of the GIN architecture.
3 Related Work
Existing post-hoc GNN explanations methods can be classified into two main categories: instance-level and model-level methods Barredo Arrieta et al. (2020). See Agarwal et al. (2023); Dai et al. (2022) for nice reviews on the subject. In the realm of instance based methods, gradient-based methods use the gradients of the output with respect to the input or intermediate features to measure the importance of each component of the graph. Decomposition-based methods try to decompose the input graph into smaller subgraphs or paths that can account for the output. Surrogate-based methods use a simpler, more interpretable model to approximate the behaviour of the original GNN and provide explanations based on the surrogate model. And finally Perturbation-based methods which perturb the input graph by removing or adding nodes, edges, or features, and observe the changes in the output to identify the influential components.
The most mainstream technique, GNNExplainer Ying et al. (2019) achieves explanation by removing redundant edges from an input graph instance, maximising the mutual information between the distribution of subgraphs and the GNN’s prediction. It is able to provide an explanation both in terms of a subgraph of the input instance to explain, and a feature mask indicating the subset of input node features which is most responsible for the GNN’s prediction.
For model-based techniques, few methods come to mind Saha et al. (2022); Azzolin et al. (2023); Vu & Thai (2020); Wang et al. (2023); Xuanyuan et al. (2023); Yuan et al. (2020); Zhang et al. (2021). The most mainstream method seems to be XGNN Yuan et al. (2020). The authors of XGNN investigate the possible input characteristics used by a GNN for graph classification. But they formulate the problem as a reinforcement learning problem and generate graph patterns iteratively. Such an iterative approach is often intractable for large graphs. Moreover, it does not allow for both node classification and graph classification explanations, nor does it allow for an investigation of the learning process through the different layers of the GNN.
4 Datasets
All three datasets have the same setup: given a set of graphs $\left\{\mathcal{G}^{1},\mathcal{G}^{2},\ldots,\mathcal{G}^{N}\right\}$, predict the corresponding binary labels$\left\{y^{1},y^{2},\ldots,y^{N}\right\}$.
The Grid-House dataset
inspired by (Agarwal et al., 2023) is designed to evaluate the compositionality of Graph Neural Network (GNN). It features two concepts: a 3x3 grid and a house-shaped graph made of five nodes. The dataset consists of Barabási-Albert (BA) graphs (Barabási, 2009) with a normal distribution of the number of nodes. The negative class includes a BA graph connected to either a grid or a house, while the positive class contains a BA graph connected to both a grid and a house (see fig. 2). In order to ensure that the average number of nodes is the same between classes, the number of nodes is a uniformly distributed between 6 and 21 for the grid graphs, between 7 and 22 for the house graphs, and between 1 and 16 when both are present. During generation, we ensure no test set leakage by removing isomorphisms. On 2,000 graphs, we perform an 80/20 train/test split.
For accurate classification, models need to identify and combine simple patterns. Recognizing isolated patterns or single node features is not sufficient. The dataset helps investigate how GNNs combine multiple concepts and addresses the “laziness” phenomenon, where networks learn patterns characterising only one class and predict the other by default (Longa et al., 2023b).
The dataset has been structured such that an optimal, linearly separable solution requires the combination of local properties, such as eigenvector centrality and betweenness centrality, or the identification of global structural motifs, like counting the number of squares (i.e., four-node cycles). A random Barabási-Albert graph can’t contain any four-node cycles, while a grid subgraph will consistently exhibit four such cycles. A house subgraph contains exactly one four-nodes cycle and one three-nodes cycle. Therefore, a graph that contains both a grid and a house will have a total of five four-node cycles. The presence of a three-node cycle could help the diagnostic of one type of graph in the negative class but is not necessary nor sufficient for solving the classification problem. On the contrary, counting the number of four-node cycle is necessary and sufficient. Thus, distinguishing between the classes does not really necessitates leveraging centrality-based measures but only recognizing the presence of a specific number of four-node cycles, enabling the model to effectively differentiate between the positive and negative classes. Thus the interesting results of fig. 3.
Figure 2: Examples of the grid-house dataset. There are graphs with only a house and only a 3x3 grid, these are in one class. Graphs with both a house and a grid are in the other class.
ClinTox Molecular
contains molecular graphs representing compounds with binary labels indicating whether they are toxic or non-toxic. The dataset consists of 1,491 drug compounds with known chemical structures. Each molecule is represented as a graph where nodes correspond to atoms and edges to bonds, with node features representing atom types and edge features representing bond types. The task is to predict toxicity.
fMRI FC connectomes
consists of two parts. The Autism Brain Imaging Data Exchange I dataset contains 528 ASD patients and 571 typically developed (TD) individuals, the REST-meta-MDD dataset contains 848 MDD patients and 794 healthy controls. For both, the task is to classify these. We use the datasets with functional connectivity (FC) graphs, as prepared by  Zheng et al. (2023). We perform a 95/5% train-test random split.
In our paradigm, we hope probing functional connectivity matrices (FC) matrices (Farahani et al., 2019) of neurological disorders (ND) could help explore the link between stuctural properties of the brain’s functional connectivity and neurological disorders such as Autism Spectrum Disorders (ASD) and Major Depressive Disorders (MDD). Similarly that probed graph properties in toxic molecules align with chemical knowledge.
5 Methodology
For each of the three datasets, we use a similar network architecture consisting of a number of GNN (GCN, GIN, or GAT) layers, followed by a pooling operation (mean- (Kipf & Welling, 2017), sum-(Xu et al., 2019), or max-pooling (Hamilton et al., 2017)), and then a number of dense layers. We optimize the hyperparameters to obtain good models for the binary classification task.
For the Grid-House dataset the hyperparameter information can be found in table 4. On this dataset we also compared different regularisation methods. The explicit $L_{2}$ regularisation encourages the network to keep the weights small, and we expect that this would make the embeddings less sensitive to fluctuations in the input data and smoother. the latter would make them more linearly separable for our probing methods. Dropout randomly disables a fraction of the neurons during each training iteration which forces the network to learn redundant representations, as any neuron could be dropped out. These redundant representations might make it more difficult to linearly separate the graph embeddings. We ran each model 20 times and took the one with the best accuracy.
For the ClinTox Molecular dataset, we ranged the number of layers from 4 to 6 and hidden dimensions from 64 to 256. The final model architectures were selected based on optimal performance on the ClinTox dataset.
For fMRI FC connectomes the hyperparamter search space is described in table 15
5.1 Probing strategy
Probing is performed on the train and test sets, where train features $\{f_{\text{train}}^{(i)}\}$ and graph properties $\{z_{\text{train}}^{(i)}\}$ are paired for each graph (equally for the test set). Let’s define at least one example for the GCN model. Let $\mathcal{G}^{i}=(A^{i},X^{i})$ denote the $i$-th graph, where $A^{i}$ is the adjacency matrix and $X^{i}$ is the node feature matrix as previously defined. The GCN layers iteratively update the node features $H^{(l)}$ through graph convolutions defined previously as $H^{(l+1)}=\sigma(\hat{A}H^{(l)}W^{(l)})$, where $\hat{A}$ is the normalized adjacency matrix, $W^{(l)}$ are the trainable weights, and $\sigma$ is a non-linear activation function (ReLU). The node embeddings $H^{(l)}$ at each layer $l$ capture both local and global structural information by aggregating features from neighboring nodes. The final node embeddings $H^{(4)}$ are pooled using global max pooling to generate a graph-level embedding $H_{\text{global}}$, which is passed through three fully connected layers to produce the final prediction $\hat{y}$. We deinfe these post pooling operations as $H_{\text{global}}^{(5)}=\sigma(W_{1}H_{\text{global}}),H_{\text{global}}^{(6)}=\sigma(W_{2}z_{1}),\hat{y}=\text{Softmax}(W_{3}z_{2})$. For probing purposes, we use $H^{(l)}$ at different layers to evaluate node-level properties, while $H_{\text{global}},H_{\text{global}}^{(5)},H_{\text{global}}^{(6)}$, and $\hat{y}$ are used to assess graph-level properties.
We aggregate node embeddings across all graphs to train a single probing classifier for each graph property. For each property, we construct a feature matrix by combining embeddings across all graphs, layer per layer. The classifier $g$ is then trained on this aggregated dataset to predict graph properties $z_{k}^{(i)}$, where $i$ denotes the $i$-th graph and $k$ represents the $k$-th graph property, as defined in table 3. This approach assumes that the relationships between node or graph embeddings and properties are consistent across graphs.
[…]
To address this, we first sort the embeddings in descending order based on their norms before concatenating, which introduces permutation invariance. Sorting in this way ensures that any padding zeros align at the end of the sequence, enabling learnable representations for graphs with varying node counts. While sorting for permutation invariance is not widely discussed in the literature, it provides a practical solution by using the embeddings’ properties to enforce consistent ordering across graphs.
6 Results
6.1 Grid-House dataset
The models performed as anticipated thanks to their high expressiveness and the linearly separable nature of the classification problem as we can see in table 5. The probing results on the Grid-House dataset demonstrate that the number of squares consistently yields the highest $R^{2}$ scores across all models in the global graph embeddings (after pooling aggregation has been applied). This aligns with our initial hypothesis.
In general, higher-layer embeddings filter out many other graph properties as they are less relevant for making the classification problem linearly separable. The GNN’s final layer focuses on the number of squares, effectively partitioning the graphs into two classes: those with #$\#\text{squares}<5$ (indicating either the grid or house alone) and those with #$\#\text{squares}=5$ (indicating the presence of both substructures). This reduction in feature space through the layers aligns with the model’s goal of optimizing the decision boundary for binary classification, where the number of squares becomes a clear and dominant factor for separability. Further confirming expectations, density and average path length are also prominent as the presence of both a house and a grid does slightly increase the average density and path length of graphs. These findings confirm the correspondence between graph embeddings clustering and property hyperplane separation as shown in fig. 5.
Figure 3: T-SNE visualisation across different layers of our GIN architecture aligned with the probing $R^{2}$ scores plots (Grid House)
We further observe that, for both the GCN and GIN models, the application of $L_{2}$ regularization yields the expected behavior. The last layer of the GCN in fig. 7 shows a stronger dominance of the number of squares feature when $L_{2}$ regularization is applied compared to when it is not. Similarly, in the GIN, both number of triangles and density become less detectable relative to the number of squares, by the probing classifier in the final layers under $L_{2}$ regularization, consistent with the anticipated effects on the feature representation.
We observe results consistent with our expectations for the models with dropout. The key property is less dominant, and multiple properties are represented in the final layers. Notably, in fig. 8 the last layer, the separability gap between the #square and the other properties is reduced, indicating a more distributed representation of features when dropout is applied.
[…]
The GAT model stands out by capturing not only squares ($R^{2}=0.88$), but also performing well on other properties like triangles, cliques, and density. GAT assigns a weight to each neighbouring node based on a learned function of the node features, aggregating the neighbours’ information in a weighted manner. This feature-dependent mechanism introduces flexibility but also makes GAT’s performance contingent on the quality and richness of the node features. It seems that GAT’s broader capability compared to the GCN comes at the cost of focus, as GAT tends to incorporate multiple features, which may dilute its ability to pinpoint the most crucial property (in this case, the number of squares) for the classification task. This over-reliance on feature aggregation can lead to inefficiencies when simpler, more targeted properties suffice, as seen with GIN. These results also make sense with regard to the best score obtained by the GIN architecture as seen in table 5.
6.2 ClinTox Molecular
As expected GIN outperformed the other models. Based on scores and good properties (inferential mechanism with better expressivity) we focus on the GIN results for the results on other datasets. Detailed results can be found in the appendix, table 8. When looking into the linear probing performance in table 1, we find that the highest scores are consistently yielded by the average degree, the spectral radius, the algebraic connectivity and the density, in that order.
Table 1: Linear Probing $R^{2}$ Performance Across GIN Layers for Selected Graph Properties (ClinTox Dataset). Best Scores in Bold; Non-convergence indicated by —
GIN Layer Avg. degree Spectral radius Alg. co. Density Avg. btw. cent. Graph energy
x_global 0.81 0.74 0.67 0.58 0.48 0.44
x6 (MLP) 0.80 0.74 0.66 0.58 0.42 0.44
x7 (MLP) 0.75 0.71 0.56 0.50 0.47 0.46
x8 (MLP) 0.07 0.02 0.00 0.06 0.05
The average degree of atoms in a molecule provides a straightforward interpretation, as atoms with higher valencies are generally less stable and less biologically compatible. For instance, hydrogen with a valency of 1 and oxygen with a valency of 2 are more compatible with carbon-based molecules, whereas sulfur, with a valency of 6, is less favorable for biological systems  (Komarnisky et al., 2003). Therefore, the average degree serves as a useful indicator of molecular toxicity. Additionally, the spectral radius, often associated with molecular stability and reactivity, is another valuable graph property. Molecules with a lower spectral radius tend to be more stable, while those with a higher spectral radius may exhibit localized electron densities, increasing their reactivity. Using this property to predict molecular toxicity is a logical approach. To the best of our knowledge, there is no comprehensive analysis exploring the role of spectral radius in the emergence of molecular toxicity, highlighting an opportunity for future research.
6.3 fMRI FC connectomes
In the detailed results, in the appendix table 14, GIN outperforms the other architectures in both parts of the dataset (reproducing the observation by Zheng et al. (2023). Again, the strength of GIN lies in its injective aggregation mechanism. The probing results on the ASD dataset reveal that the number of triangles consistently achieves high $R^{2}$ scores across all models, with particularly strong performance in GIN models. This property is followed by the spectral radius and the density.
As further detailed in the appendix tables 16, 18, 19 and 17 the number of edges is particularly well encoded in the representation of the GAT. This is a consequence of its reliance on feature-dependent weights and structure-free normalisation, which limit its ability to capture specific structural properties that do not directly depend on edges. The GCN results are broadly comparable to those of the GIN, though they tend to be less precise and selective.
For the MDD results we also focus the GIN model. Detailed results are in the appendix tables 20, 23, 24 and 22. The probing results MDD reveal that the number of triangles still consistently achieves high $R^{2}$ scores across all models while being less of a distinctive feature than in ASD. This time, the spectral radius is dominated by the density of the graph. In general, the embeddings from the 7th layer of our GIN architecture exhibit higher $R^{2}$ scores for relevant graph properties, suggesting improved separability in the embedding space for MDD classification compared to ASD. This indicates that the learned representations at this depth capture more discriminative structural features, facilitating more effective class separation between MDD and healthy controls.
7 Discussion
7.1 Expectations
For Grid-House we hypothesized that the GNN would benefit from leveraging both the local clustering coefficient and eigenvector centrality as node-level features. The first one would help characterize a house, the second a grid. However, neither feature alone is sufficient to render the problem linearly separable. We therefore expect either a combination of features or a single global property (e.g. number of squares) to be leveraged. If the tensor embeddings produced by the GNN can be used to predict these properties, this would indicate that the GNN is utilising them in solving the classification task.
For the ClinTox Molecular dataset, based on the literature Kengkanna & Ohue (2024); Chen et al. (2021); Jiang et al. (2021) some few properties have been found to be link with toxicity such as the node degree (i.e. the valency), subgraph patterns (functional groups, chemical fragments), and the overall graph connectivity.
Based on existing literature on functional connectivity (FC) network properties in ASD and MDD, we hypothesized that specific properties will be critical in classifying brain networks for the fMRI FC connectomes dataset. For ASD, we expect betweenness centrality to play a significant role at the node level, reflecting local overconnectivity. At the graph level, we anticipate that clustering coefficient, characteristic path length, and small-worldness will be essential in capturing the local and global network disruptions seen in ASD, particularly the imbalance between local overconnectivity and long-range underconnectivity. For MDD, we hypothesise that increased clustering coefficients, modularity, number of triangles and number of squares will be key features for classification, as they could indicate of heightened local interconnectedness and disrupted global integration.
7.2 Findings
We first demonstrate the feasibility of our probing method through the Grid-House dataset. This acts as a proof of concept on probing classifiers plumbing the representations learned by the GNN. We made sure to choose a classification task which requires learning global structural properties of the graph, such as motifs like squares and triangles, and a long range dependencies between those. In line with our expectations, the results show significant dominance of number of squares in the post pooling layers of every model, while still highlighting the superiority of the GIN model in leveraging superior representations when trained to classify graphs with complex motifs like Grid and House. This shows the expressivity of the models and their ability to reduce the complexity of a graph related problem to known graph properties, making the problem linearly separable in the space of their embeddings. These results are consistently higher than those obtained from probing the models with randomized labeling, highlighting the relevance of the initial findings.
Using the ClinTox Molecular dataset to assess molecular toxicity, we explored how key graph properties, such as the average degree and spectral radius, are utilized by our GIN architecture. The average degree, closely linked to atomic valency, reflects a molecule’s potential for interactions. The spectral radius offers a complementary hypothesis, suggesting that the overall structural stability of a molecule, independent of specific atomic features, may also be a key factor in toxicity prediction. These results suggest that despite the limitations of our approach, it still holds potential for assisting research in complex systems fields, such as neuroscience or social sciences, where emergent phenomena play a crucial role in understanding system dynamics.
Given the previous positive results, we explored a real life applications with the fMRI FC connectomes dataset. Here, the results provide new insights that extended beyond our initial hypothesis. While we expected betweenness centrality, clustering coefficient, and characteristic path length and Small-worldness to be the most relevant for distinguishing ASD from healthy individuals, the prominence of the number of triangles highlighted the importance of local structural motifs. This makes sense in the context of functional connectivity, where local overconnectivity in specific brain regions, such as sensory and association cortices, has been observed in individuals with ASD. The strong role of triangle motifs may reflect the tight, redundant local connections that characterize these regions, supporting the hypothesis that local overconnectivity is a key factor in ASD. The spectral radius and density and graph energy being particularly significant is also logical, as these properties are closely related to the overall connectivity strength and the compactness of connections within subnetworks. The presence of SW is still quite significant in the post-pooling layers which is interesting. In ASD, where global integration is often reduced and local connectivity heightened, these metrics may provide an important reflection of the imbalance between short-range and long-range communication pathways in the brain.
The outcomes for both the ASD and the MDD datasets showed promising results that should be discussed more deeply with neuroscientists. The results mainly suggest the importance of graph substructures or spectral and small-world properties over more basic graph properties to explain how graph neural networks predict these neurological disorders in the FC matrices of patients. The presence of algebraic connectivity in the last two layers of the GIN MDD probing setting while being completely absent from the GIN ASD probing one makes it a property of choice to examine for understanding better how this kind of neurological disorders affect the brain’s connectivity. This is a result which, to the best of my knowledge, has never been investigated before.
7.3 Comparison between datasets
Comparing the ClinTox and the fMRI datasets an interesting observation emerges: basic graph properties (such as the number of nodes, of edges or the average path length) are almost omnipresent in the early layers of the GIN trained on the ClinTox dataset. However, their presence is less pronounced in the GIN trained on the ASD or MDD datasets. This difference offers a clue in distinguishing the complexity of brain-related neurological disorders from the complexity of chemical qualities such as toxicity. This suggests that the emergent properties of the brain may not be as easily tied to simple, differentiable structural features as those seen in molecular systems.
As a confirmation, the types of global graph properties present in the post pooling layers of the GIN-clintox model are of less high level of abstraction than the ones in GIN-MDD or GIN-ASD. The presence of the average degree, the spectral radius, the algebraic connectivity and the density as accurate explanations for the prediction of toxicity in molecules. The presence of the spectral radius in the last layer of the GIN makes it an even more interesting property to study for toxicity. On the other hand, the presence of motifs should be more investigated in the ASD and MDD datasets with eventually more complex motifs being probed (hexagons constituted of neighbored triangles, house, grid, etc).
7.4 Future work
Our methodology has several limitations. While we addressed dataset issues such as leakage and isomorphic graphs, a key challenge remains the lack of guarantees that GNNs find globally optimal solutions, despite their theoretical capacity as universal function approximators. This is particularly evident in fMRI data, where multiple layers of complexity—from MRI limitations and BOLD signal characteristics to Pearson correlation for functional connectivity—introduce noise and inaccuracies. Investigating additional graph properties like girth or complex motifs could be beneficial. Preliminary work on alternative architectures (e.g., GATv2, GraphSAGE, ChebNet, Set2Set, HO-Conv, DiffPool) has begun but is not yet complete.
8 Conclusion
We demonstrate the relevance of our model-agnostic explainability method for graph neural networks which probe for GNN graph theoretic representations on the Grid-House dataset. We anticipate any lazy learning bias. We manifest both the expressivity of different GNN architectures and their ability to solve a graph classification problem through optimal feature extraction. They render it linearly separable in the space of their embeddings through the computation of the number of squares in the graph.
That experiment prompted us to investigate both the Clintox Molecular dataset and fMRI FC connectomes dataset and anchored the possibility of formulating hypotheses on the emergent dependence of complex systems qualities to basic and more higher level structural properties. This kind of higher order properties are very interesting in order to understand how the macroscopic behavior of complex systems emerges from the intricate interplay of their microscopic components. For example how diseases spread in social networks, how neurons interact in the brain, or how information propagates through the Internet. There is a manifest emergence of molecular qualities like toxicity with regard to their structural properties like node degree (atom valency) and spectral radius (the molecule’s stability). But the complexity of complex systems like the brain makes blurrier the possibility of understanding what affects what as, for example, one could argue that behavioural therapies might influence the brain connectivity as the brain connectivity might influence behavioural qualities of one patient. Echoing this egg-chicken conundrum, the investigation of motifs and spectral properties’ role in neurological disorders like ASD and MDD could allow for promising avenues.
Author Contributions
Pelletreau-Duris Tom: Conceptualization (lead); methodology (lead); software (lead); formal analysis (lead); writing – original draft (lead); writing – review and editing (equal). Ruud van Bakel: writing – review and editing (equal); Supervision Michael Cochez:Writing – review and editing (equal); Supervision; Funding acquisition
Acknowledgments
RvB and MC are in part funded by the Graph-Massivizer project, the Horizon Europe research and innovation program of the European Union grant management number 101093202 https://graph-massivizer.eu/
Appendix A R2 Score
[…]
From an information-theoretic perspective, training the probing classifier $g$ can be viewed as estimating the mutual information between the learned representations $f_{l}(x)$ and the property $z$. This mutual information is denoted as $\mathrm{I}(\mathbf{z};\mathbf{h})$, where $\mathbf{z}$ refers to the property and $\mathbf{h}$ represents the intermediate representations Belinkov (2021).
Appendix B Local and global graph properties
Local Property Visual Pattern & Definition Computational Criteria
Degree How many links a node has which is the simplest form of centrality Count edges per node
Local clustering Coefficient Are the neighbours of a node also connected together ? Count triangles of neighbours / total possible triangles of neighbours
Betweenness Centrality How much of a bridge between clusters is a node. Removing that node would break many shortest paths. Importance in information flow Number of shortest paths through node
Closeness Centrality Being in the middle of the network, the barycenter of the graph. The average length of the geodesic distances to all the other nodes (inverse sum of shortest paths)
Eigenvector Centrality Being connected to well connected nodes without necessarily having a large number of neighbours itself; influence based on connections Recursive definition based on neighbours
PageRank Nodes with important connections; web-inspired importance Similar to Eigenvector but with random walk and teleportation
Table 2: Local Network Properties with definition and computational criteria
Global Property Visual Pattern & Definition Computational Criteria
Number of Nodes Graph size; total nodes in the network Count vertices
Number of Edges Graph density; total connections in the network Count connections
Density Overall graph connectivity; how densely connected Ratio of actual to possible edges
Average Path Length On average, how close are nodes to each other? Typical distance between node pairs Average number of steps along the shortest paths for all possible pairs of nodes
Diameter Graph span; longest of all shortest paths Maximum shortest path
Radius Graph core; minimum distance from central to farthest node Minimum eccentricity
Transitivity Triangle density; probability of connected node triplets Ratio of triangles to triads
Assortativity Node degree correlations; tendency of similar nodes to connect Pearson correlation of degrees
Number of Cliques Dense subgraphs; count of maximal fully connected subgraphs Number of maximal complete subgraphs
Number of Triangles Local density; fully connected 3-node subgraphs Count 3-node cliques
Number of Squares 4-node patterns; cycles in the graph Count 4-node cycles
Largest Component Size Main connected structure; size of biggest connected part Largest set of connected nodes
Average Degree Overall connectivity; average connections per node Mean of all node degrees
Spectral Radius Dominant graph structure; overall connectivity measure Largest eigenvalue of adjacency matrix
Algebraic Connectivity Graph cohesion; measure of how well-connected the graph is Second smallest eigenvalue of Laplacian
Graph Energy The eigenvalues capture deviations from regularity in the network. Complete graphs or highly connected networks tend to have higher energies due to the larger magnitude of their eigenvalues. In social networks, biology, and communication networks, graph energy can help assess robustness, synchronizability. Sum of absolute Laplacian eigenvalues
Small World Coefficient Balance of clustering and paths; small-world characteristics Comparison to random graph
Small World Index Refined small-world measure; comparison to random and lattice graphs Comparison to random and lattice graphs
Betweenness Centralization Central node dominance; degree of central bridging node Variation in betweenness centrality across nodes
PageRank Centralization Influence concentration; degree of dominant influential nodes Variation in PageRank values across nodes
Table 3: Global Network Properties with definition and computational criteria
B.1 Artificial dataset
B.1.1 Grid House figures
Figure 4: Comparison of different centrality measures for the first graph in our Grid House dataset: (a) betweenness centrality, (b) eigenvector (PageRank) centrality, and (c) local clustering coefficients.
B.1.2 Grid House models
Table 4: Range of Hyper-parameters and Final Specification for the Grid-House Dataset
Hyper-parameter Range Examined Final Specification
Graph Encoder
#GNN Layers {$[2,3,4,5]$} 4 (GCN), 2 (GIN), 3 (GAT)
#MLP Layers {$[2,3,4]$} 3 (GCN), 2 (GIN), 2 (GAT)
Hidden Dimensions {$[10,15,30,45,60,64,128,256]$} 60 (GCN), 30 (GIN), 128 (GAT)
Attention Heads (GAT) {$[4,8,16]$} 8 heads, 32 dimensions per head
Learning Rate {$[1\mathrm{e}-2,1\mathrm{e}-3,1\mathrm{e}-4]$} $31\mathrm{e}-$ 3
Batch Size {$[32,64,128,256]$} 64
Weight Decay (when added) {$[1\mathrm{e}-4,1\mathrm{e}-2]$} $41\mathrm{e}-$ 4 (GCN), $21\mathrm{e}-$ 2 (GIN)
Batch Normalization {$with,without$} $without$
Dropout (when added) {[0.15, 0.5]} 0.2
Pooling Method {$mean,sum,max$} $max$ (GCN), $mean$ (GIN), $max$ (GAT)
Table 5: Performance of Different Models with Regularization on the Artificial Dataset (80%-20% Random Split). The highest performance is highlighted with boldface. All performances are reported under their best settings and rounded to 2 decimal places.
Method Test Accuracy
GCN (control) $0.90$
GCN ($L_{2}$) $0.97$
GCN (dropout) $0.93$
GIN (control) $\mathbf{1.00}$
GIN ($L_{2}$) $0.99$
GIN (dropout) $1.00$
GAT $0.97$
As expected the RGCN outperform the GCN on this node classification task.
B.1.3 Grid House RESULTS
B.1.4 Graph properties probing results
Table 6: Linear Probing $R^{2}$ Performance Across models for Selected Graph Properties (GridHouse Dataset). Best Scores in Bold; Non-convergence indicated by —
Model #nodes #edges density avg path len #cliques #triangles #squares #Largest Component
GCN (control)
x_global 0.36 0.66 0.33 0.02 0.31 0.77 0.36
x5 0.33 0.22 0.64 0.29 0.27 0.39 0.77 0.33
x6 0.19 0.08 0.56 0.07 0.06 0.74 0.19
x7 0.45 0.13 0.03 0.72
GCN ($L_{2}$)
x_global 0.36 0.09 0.67 0.35 0.20 0.68 0.86 0.36
x5 0.31 0.32 0.66 0.32 0.32 0.80 0.86 0.31
x6 0.04 0.41 0.15 0.03 0.23 0.83 0.04
x7 0.29 0.27 0.09 0.81
GCN (dropout)
x_global 0.21 0.07 0.67 0.33 0.07 0.63 0.72 0.22
x5 0.59 0.26 0.66 0.74
x6 0.42 0.21 0.49 0.65
x7 0.35 0.10 0.26 0.51
GIN (control)
x_global 0.12 0.07 0.50 0.32 0.07 0.22 0.87 0.12
x5 0.72 0.30 0.89 0.93
x6 0.02 0.11 0.88
GIN ($L_{2}$)
x_global 0.49 0.30 0.18 0.85
x5 0.51 0.15 0.52 0.89
x6 0.40 0.12 0.10 0.80
GIN (dropout)
x_global 0.53 0.36 0.25 0.87
x5 0.71 0.33 0.85 0.93
x6 0.21 0.34 0.91
GAT
x_global 0.54 0.59 0.49 0.61 0.89 0.87 0.54
x5 0.33 0.27 0.17 0.64
x6 0.25 0.17 0.17 0.63
B.1.5 Graph properties probing plots
GCN
Figure 5: T-SNE visualisation across different layers of our GCN architecture aligned with the probing $R^{2}$ scores plots (Grid House)
Figure 6: Plot of the GCN (control) $R^{2}$ results across different layers probing for graph properties with post pooling layers only (Grid House)
Figure 7: Plot of the GCN ($L_{2}$) $R^{2}$ results across different layers probing for graph properties with post pooling layers only(Grid House)
Figure 8: Plot of the GCN (dropout) $R^{2}$ results across different layers probing for graph properties with post pooling layers only(Grid House)
GIN
Figure 9: Plot of the GIN (control) $R^{2}$ results across different layers probing for graph properties with post pooling layers only (Grid House)
Figure 10: Plot of the GIN ($L_{2}$) $R^{2}$ results across different layers probing for graph properties with post pooling layers only (Grid House)
Figure 11: Plot of the GIN (dropout) $R^{2}$ results across different layers probing for graph properties with post pooling layers only (Grid House)
GAT
Figure 12: Plot of the GAT $R^{2}$ results across different layers probing for graph properties with post pooling layers only (Grid House)
B.1.6 Grid House Node properties probing results
Using the probing method developed in the next section, we were not fully able to confirm our initial hypothesis.
Table 7: Linear Probing $R^{2}$ Performance Across models for Selected Node Properties (GridHouse Dataset). Best Scores in Bold; Non-convergence indicated by —
GCN Layer degree closeness betweenness eigenvector clustering pagerank
x1 (GCN) 0.50 0.22 0.25 0.19 0.06 0.56
x2 (GCN) 0.54 0.32 0.28 0.24 0.09 0.57
x3 (GCN) 0.54 0.35 0.29 0.25 0.11 0.57
x4 (GCN) 0.55 0.37 0.28 0.30 0.17 0.57
GIN Layer
x1 (GIN) 0.55 0.18 0.24 0.22 0.05 0.56
x2 (GIN) 0.52 0.34 0.27 0.25 0.07 0.54
GAT Layer
Layer 0 0.55 0.07 0.05 0.32 0.28 0.17
Layer 1 0.52 0.48 0.08 0.31 0.30 0.14
Layer 2 0.47 0.55 0.29 0.29
Layer 3 0.41 0.14 0.19 0.26
Layer 4 0.35 0.50 0.12 0.21 0.23
In these pre-pooling layers, we first observe the predominance of page rank and node degree in the early layers and in all the layers of the GCN and the GIN (which has only two of them). When considering the last layers of the GAT (unfortunately we should have have similar architecture with the GIN in order to fully test our hypothesis) it seems that closeness, node degree and clustering coefficient are the most significant. This aligns with our framing of the graph classification task, which is largely driven by the detection of squares and the fact that pre-pooling layers leading to this property detection should affect mostly these three properties. But this does not align with the use of node properties in a graph in order to do graph classification. This still makes a lot of sense. In general, contrary to the graph probing, and to the exception of the node degree, we see that there is not a single property clearly dominating others but that we go towards a combination of different properties just before the graph pooling method. We would have expect the GIN architecture to show similar results with four layers (as we already see an important increase with regard to the closeness between the first and second layer).
B.2 Clintox dataset
B.2.1 model
Table 8: Performance of Different Models on ClinTox with a 80%-20% Random Split. The highest performance is highlighted with boldface. All the performance of methods are reported under their best settings.
Method ClinTox
GCN $0.91$
GAT $0.92$
GIN $0.93$
B.2.2 Results
B.2.3 Graphs properties probing results
Table 9: Linear Probing $R^{2}$ Performance across the GIN layers for basic graph properties (ClinTox dataset). Best Scores in Bold; Non-convergence indicated by —(full)
GIN Layer # Nodes # Edges Density Avg. Path Length Diameter Radius
x1 (GIN) 1.00 1.00 0.66 0.76 0.55 0.60
x2 (GIN) 1.00 1.00 0.57 0.95 0.88** 0.84
x3 (GIN) 1.00 1.00 0.62 0.97 0.93 0.89
x4 (GIN) 0.99 0.99 0.37 0.91 0.82 0.82
x5 (GIN) 0.99 0.99 0.29 0.90 0.82 0.82
x_global 0.41 0.44 0.58 0.20 0.20 0.20
x6 (MLP) 0.40 0.44 0.58 0.19 0.19 0.19
x7 (MLP) 0.42 0.46 0.50 0.27 0.23 0.25
x8 (MLP) 0.04 0.05 0.00 0.04 0.05 0.03
Table 10: Linear Probing $R^{2}$ Performance across the GIN layers for clustering and centrality measures (ClinTox dataset). Best Scores in Bold; Non-convergence indicated by —(full)
GIN Layer Clustering coef. Transitivity Assortativity Avg. clustering Avg. btw. cent. PageRank cent.
x1 (GIN) 0.32 0.18
x2 (GIN) 0.21
x3 (GIN)
x4 (GIN)
x5 (GIN)
x_global 0.25 0.48 0.40
x6 (MLP) 0.27 0.42 0.39
x7 (MLP) 0.47
x8 (MLP) 0.06
Table 11: Linear Probing $R^{2}$ Performance across the GIN layers for graph substructures (ClinTox dataset). Best Scores in Bold; Non-convergence indicated by —(full)
GIN Layer # Cliques # Triangles # Squares Largest comp. size Avg. degree Graph energy
x1 (GIN) 0.99 0.00 0.99 0.53 1.00
x2 (GIN) 1.00 0.00 0.99 0.46 1.00
x3 (GIN) 1.00 0.00 0.99 0.53 1.00
x4 (GIN) 0.99 0.00 0.99 0.20 0.99
x5 (GIN) 0.99 0.00 0.99 0.99
x_global 0.43 0.00 0.40 0.81 0.44
x6 (MLP) 0.43 0.00 0.40 0.80 0.44
x7 (MLP) 0.46 0.00 0.42 0.75 0.46
x8 (MLP) 0.04 0.00 0.04 0.05
Table 12: Linear Probing $R^{2}$ Performance across the GIN layers for spectral and small-world properties (ClinTox dataset). Best Scores in Bold; Non-convergence indicated by —(full)
GIN Layer Spectral rad. Algebraic co. Small world coef. Small world idx Avg. btw. cent.
x1 (GIN) 0.70 0.78
x2 (GIN) 0.66 0.80
x3 (GIN) 0.61 0.80
x4 (GIN) 0.16 0.78
x5 (GIN) 0.69
x_global 0.74 0.67 0.48
x6 (MLP) 0.74 0.66 0.42
x7 (MLP) 0.71 0.56 0.47
x8 (MLP) 0.07 0.02 0.06
B.2.4 Plots
Figure 13: Plot of the GIN $R^{2}$ results across different layers probing for graph properties. ClinTox dataset (the negative $R^{2}$ values have been reduced to -0.05).
B.2.5 Node properties probing results
Table 13: Linear Probing $R^{2}$ Performance across the GIN layers for various node properties (ClinTox dataset). Best Scores in Bold; Non-convergence indicated by —
GIN Layer degree closeness betweenness eigenvector clustering pagerank
x0 (GIN) 0.99 0.06 0.57 0.30 0.16
x1 (GIN) 0.85 0.12 0.51 0.31 0.00 0.20
x2 (GIN) 0.89 0.11 0.59 0.29 0.26
x3 (GIN) 0.86 0.07 0.51 0.28 0.17
x4 (GIN) 0.85 0.09 0.49 0.32 0.14
Here again, the very strong presence of the node degree makes a lot of sense when we know this property prepares the aggregation of global properties in the post pooling layers. The interesting thing is the non negligible presence of the betweenness centrality in all the layers which suggests that the betweenness centrality of atoms is important in the aggregation of global molecule properties that help predict the toxicity of a molecule. This property is more than the closeness or the clustering coefficient. The irreplaceable nature of some atoms in the molecular graph, which is literally the meaning of having a high betweenness centrality, is an important feature which makes these atoms targets to be part of higher order molecular schemes and patterns.
Figure 14: Plot of the GIN $R^{2}$ results across different layers probing for node properties. ClinTox dataset (the negative $R^{2}$ values have been reduced to -0.05). (full results)
B.3 fMRI datasets
B.3.1 models
Table 14: Performance of Different Models on REST-meta-MDD and ABIDE with a 95%-5% Random Split. The highest performance is highlighted with boldface. All the performance of methods are reported under their best settings and round to the second decimal.
Method ABIDE REST-meta-MDD
GCN $0.56$ $0.61$
GIN $0.69$ $0.69$
GAT $0.62$ $0.67$
Table 15: Range of Hyper-parameters and Final Specification for FC datasets
Hyper-parameter Range Examined Final Specification
Graph Encoder
#GNN Layers {$[4,5,6]$} 5
#GIN Layers {$[4,5,6]$} 5
#GAT Layers {$[4,5,6]$} 5
#MLP Layers (for all models) {$[2,3,4]$} 2
#GCN Hidden Dimensions {$[64,128,256]$} 128
#GIN Hidden Dimensions {$[64,128,256]$} 128
#GAT Hidden Dimensions {$[64,128,256]$} 128
#GCN aggregation method {$[mean,sum,max(pooling)]$} max pooling
#GIN aggregation method {$[mean,sum,max(pooling)]$} mean pooling
#GAT aggregation method {$[mean,sum,max(pooling)]$} max pooling
GCN Learning Rate {$[1\mathrm{e}-2,1\mathrm{e}-3,5\mathrm{e}-4,1\mathrm{e}-4]$} $45\mathrm{e}-$ 4
GIN Learning Rate {$[1\mathrm{e}-2,1\mathrm{e}-3,5\mathrm{e}-4,1\mathrm{e}-4]$} $45\mathrm{e}-$ 4
GAT Learning Rate {$[1\mathrm{e}-2,1\mathrm{e}-3,5\mathrm{e}-4,1\mathrm{e}-4]$} $21\mathrm{e}-$ 2
Batch Size (all models) {$[32,64,128]$} 32
Weight Decay (alll models) {$[1\mathrm{e}-3,1\mathrm{e}-4]$} $41\mathrm{e}-$ 4
batch normalisation {$with,without$} $without$
dropout {$with,without$} $without$
B.3.2 RESULTS ABIDE (ASD) dataset
Table 16: Linear Probing $R^{2}$ Performance across GNN layers for basic graph properties (ASD dataset). Best Scores in Bold; Non-convergence indicated by —(full)
GCN Layer # Nodes # Edges Density Avg. Path Length Diameter Radius
x1 (GCN) 0.90 0.21 0.13 0.07
x2 (GCN) 0.77 0.22 0.24
x3 (GCN) 0.62 0.31
x4 (GCN) 0.38 0.14
x5 (GCN) 0.02 0.09
x_global 0.58 0.56 0.48 0.36 0.37
x6 (MLP) 0.52 0.50 0.45 0.39 0.41
x7 (MLP)
GIN Layer
x1 (GIN) 0.94 0.41 0.47 0.45
x2 (GIN) 0.55 0.38 0.28 0.23
x3 (GIN) 0.25 0.25
x4 (GIN)
x5 (GIN) 0.18
x_global 0.56 0.58 0.11 0.07 0.00
x6 (MLP) 0.58 0.66 0.14 0.10 0.09
x7 (MLP) 0.36 0.37 0.09 0.11
x8 (MLP)
GAT Layer
x (GAT) 0.93 0.16 0.04
x2 (GAT) 0.89 0.16 0.34 0.29
x3 (GAT) 0.84 0.30 0.39 0.31
x4 (GAT) 0.78 0.27 0.48 0.08
x5 (GAT) 0.67 0.52 0.44
x_global 0.74 0.70 0.60 0.29 0.40
x6 (GAT) 0.82 0.81 0.56 0.46 0.48
x7 (GAT)
Table 17: Linear probing performance (R2 score) across GCN layers for clustering and centrality measures (ASD dataset). Best Scores in Bold; Non-convergence indicated by —(full)
GCN Layer Clustering coe. Transitivity Assortativity Avg. clustering Avg. btw. cent. PageRank cent.
x1 (GCN)
x2 (GCN)
x3 (GCN)
x4 (GCN)
x5 (GCN)
x_global 0.48 0.52 0.05 0.48 0.45 0.14
x6 (MLP) 0.33 0.30 0.33 0.41 0.06
x7 (MLP)
GIN Layer
x1 (GIN)
x2 (GIN)
x3 (GIN)
x4 (GIN)
x5 (GIN)
x_global 0.19 0.04 0.19 0.12
x6 (MLP) 0.23 0.08 0.23
x7 (MLP) 0.04 0.09 0.11
x8 (MLP)
GAT Layer
x (GAT)
x2 (GAT)
x3 (GAT) 0.02 0.02
x4 (GAT)
x5 (GAT)
x_global 0.44 0.08 0.41
x6 (GAT) 0.53 0.49 0.01 0.53 0.08
x7 (GAT) 0.00 0.00
Table 18: Linear probing performance (R2 score) across GCN layers for graph substructures (ASD dataset). Best Scores in Bold; Non-convergence indicated by —(full)
GCN Layer # Cliques # Triangles # Squares Largest comp. size Avg. degree Graph energy
x1 (GCN) 0.51 0.88 0.54 0.85 0.90
x2 (GCN) 0.27 0.81 0.58 0.77 0.77
x3 (GCN) 0.06 0.73 0.40 0.64 0.62
x4 (GCN) 0.64 0.11 0.30 0.39
x5 (GCN) 0.61 0.04
x_global 0.46 0.42 0.61 0.19 0.57 0.58
x6 (MLP) 0.42 0.34 0.35 0.31 0.51 0.52
x7 (MLP) 0.00
GIN Layer
x1 (GIN) 0.58 0.95 0.69 0.94 0.95
x2 (GIN) 0.91 0.12 0.64 0.56
x3 (GIN) 0.74 0.22 0.25
x4 (GIN) 0.54
x5 (GIN) 0.75 0.23 0.17
x_global 0.86 0.14 0.57 0.56
x6 (MLP) 0.86 0.12 0.54 0.60
x7 (MLP) 0.59 0.00 0.37 0.36
x8 (MLP)
GAT Layer
x (GAT) 0.58 0.86 0.66 0.93 0.93
x2 (GAT) 0.56 0.82 0.69 0.87 0.89
x3 (GAT) 0.54 0.80 0.70 0.80 0.84
x4 (GAT) 0.50 0.75 0.74 0.75 0.78
x5 (GAT) 0.24 0.72 0.59 0.71 0.67
x_global 0.32 0.56 0.40 0.73 0.74
x6 (GAT) 0.51 0.76 0.52 0.20 0.81 0.82
x7 (GAT)
Table 19: Linear probing performance (R2 score) across GCN layers for spectral and small-world properties (ASD dataset). Best Scores in Bold; Non-convergence indicated by —(full)
GCN Layer Spectral rad. Algebraic co. Small world coe. Small world idx Avg. btw. cent.
x1 (GCN) 0.72
x2 (GCN) 0.74
x3 (GCN) 0.56
x4 (GCN) 0.36
x5 (GCN)
x_global 0.46 0.43 0.48 0.45
x6 (MLP) 0.38 0.41 0.39 0.41
x7 (MLP) 0.00
GIN Layer
x1 (GIN) 0.88
x2 (GIN) 0.43
x3 (GIN) 0.25
x4 (GIN)
x5 (GIN) 0.24
x_global 0.76 0.40 0.12
x6 (MLP) 0.74 0.41
x7 (MLP) 0.18 0.23 0.11
x8 (MLP)
GAT Layer
x (GAT) 0.79
x2 (GAT) 0.77
x3 (GAT) 0.02 0.02
x4 (GAT) 0.64
x5 (GAT) 0.49 0.09
x_global 0.58 0.20 0.38 0.56
x6 (GAT) 0.74 0.56 0.16 0.62 0.54
x7 (GAT) 0.00
B.3.3 RESULTS REST-meta-MDD (MDD) dataset
Table 20: Linear probing performance (R2 score) across GNN layers for basic graph properties (MDD dataset). Best Scores in Bold; Non-convergence indicated by —(full)
GCN Layer # Nodes # Edges Density Avg. Path Length Diameter Radius
x1 (GCN) 0.90
x2 (GCN) 0.85
x3 (GCN) 0.71
x4 (GCN) 0.64
x5 (GCN) 0.03
x_global 0.63 0.76 0.70 0.47 0.32 0.29
x6 (MLP) 0.60 0.67 0.60 0.33 0.23 0.18
x7 (MLP)
GIN Layer
x1 (GIN) 0.85 0.50
x2 (GIN) 0.67
x3 (GIN)
x4 (GIN)
x5 (GIN)
x_global 0.55 0.89
x6 (MLP) 0.55 0.60
x7 (MLP) 0.74 0.77
x8 (MLP)
GAT Layer
x (GAT) 0.94 0.04
x2 (GAT) 0.91
x3 (GAT) 0.86
x4 (GAT) 0.84
x5 (GAT) 0.73 0.20 0.16
x_global 0.52 0.80 0.74 0.29
x6 (GAT) 0.62 0.76 0.69 0.43 0.18 0.26
x7 (GAT)
Table 21: Linear probing performance (R2 score) across GCN layers for clustering and centrality measures (MDD dataset). Best Scores in Bold; Non-convergence indicated by —(full)
GCN Layer Clustering coe. Transitivity Assortativity Avg. clustering Avg. btw. cent. PageRank cent.
x1 (GCN)
x2 (GCN)
x3 (GCN)
x4 (GCN)
x5 (GCN)
x_global 0.42 0.34 0.42 0.33
x6 (MLP) 0.35 0.33 0.35 0.41 0.11
x7 (MLP)
GIN Layer
x1 (GIN)
x2 (GIN)
x3 (GIN)
x4 (GIN)
x5 (GIN)
x_global
x6 (MLP) 0.22 0.22
x7 (MLP) 0.43 0.33 0.43
x8 (MLP)
GAT Layer
x (GAT)
x2 (GAT)
x3 (GAT) 0.02
x4 (GAT)
x5 (GAT)
x_global 0.45 0.59 0.45 0.24
x6 (GAT) 0.53 0.44 0.53 0.16
x7 (GAT)
Table 22: Linear probing performance (R2 score) across GCN layers for clustering and centrality measures (MDD dataset). Best Scores in Bold; Non-convergence indicated by —(full)
GCN Layer Clustering coe. Transitivity Assortativity Avg. clustering Avg. btw. cent. PageRank cent.
x1 (GCN)
x2 (GCN)
x3 (GCN)
x4 (GCN)
x5 (GCN)
x_global 0.42 0.34 0.42 0.33
x6 (MLP) 0.35 0.33 0.35 0.41 0.11
x7 (MLP)
GIN Layer
x1 (GIN)
x2 (GIN)
x3 (GIN)
x4 (GIN)
x5 (GIN)
x_global
x6 (MLP) 0.22 0.22
x7 (MLP) 0.43 0.33 0.43
x8 (MLP)
GAT Layer
x (GAT)
x2 (GAT)
x3 (GAT) 0.02
x4 (GAT)
x5 (GAT)
x_global 0.45 0.59 0.45 0.24
x6 (GAT) 0.53 0.44 0.53 0.16
x7 (GAT)
Table 23: Linear probing performance (R2 score) across GNN layers for graph substructures (MDD dataset). Best Scores in Bold; Non-convergence indicated by —(full)
GCN Layer # Cliques # Triangles # Squares Largest comp. size Avg. degree Graph energy
x1 (GCN) 0.52 0.77 0.57 0.88 0.90
x2 (GCN) 0.58 0.84 0.69 0.83 0.85
x3 (GCN) 0.26 0.80 0.55 0.72 0.72
x4 (GCN) 0.04 0.79 0.51 0.52 0.64
x5 (GCN) 0.52 0.01 0.04
x_global 0.54 0.76 0.50 0.62 0.73 0.76
x6 (MLP) 0.55 0.66 0.44 0.62 0.63 0.67
x7 (MLP) 0.06 0.10 0.08 0.09
GIN Layer
x1 (GIN) 0.09 0.98 0.58 0.86 0.85
x2 (GIN) 0.97 0.45 0.48 0.67
x3 (GIN) 0.87 0.05
x4 (GIN) 0.65
x5 (GIN) 0.22
x_global 0.91 0.70 0.58
x6 (MLP) 0.85 0.67 0.54
x7 (MLP) 0.02 0.88 0.51 0.75 0.74
x8 (MLP) 0.02
GAT Layer
x (GAT) 0.67 0.82 0.70 0.07 0.93 0.94
x2 (GAT) 0.59 0.83 0.81 0.89 0.91
x3 (GAT) 0.56 0.82 0.83 0.82 0.86
x4 (GAT) 0.51 0.87 0.79 0.82 0.84
x5 (GAT) 0.45 0.83 0.26 0.67 0.74
x_global 0.56 0.79 0.68 0.56 0.78 0.80
x6 (GAT) 0.53 0.76 0.55 0.65 0.73 0.76
x7 (GAT)
Table 24: Linear probing performance (R2 score) across GNN layers for spectral and small-world properties (MDD dataset). Best Scores in Bold; Non-convergence indicated by —(full)
GCN Layer Spectral rad. Algebraic co. Small world coe. Small world idx Avg. btw. cent.
x1 (GCN) 0.52
x2 (GCN) 0.60 0.20
x3 (GCN) 0.53
x4 (GCN) 0.47
x5 (GCN) 0.07
x_global 0.60 0.63 0.28 0.31 0.33
x6 (MLP) 0.51 0.58 0.23 0.41 0.16
x7 (MLP) 0.04 0.04 0.00 0.00
GIN Layer
x1 (GIN) 0.64
x2 (GIN) 0.52
x3 (GIN) 0.64
x4 (GIN)
x5 (GIN)
x_global 0.75 0.32 0.44 0.39
x6 (MLP) 0.73 0.20 0.43 0.40
x7 (MLP) 0.70 0.60 0.30 0.36
x8 (MLP) 0.03 0.01 0.00 0.01
GAT Layer
x (GAT) 0.70 0.02 0.00
x2 (GAT) 0.66 0.23
x3 (GAT) 0.68 0.26
x4 (GAT) 0.73 0.30
x5 (GAT) 0.66 0.12
x_global 0.68 0.63 0.18 0.52 0.04
x6 (GAT) 0.63 0.59 0.21 0.50 0.25
x7 (GAT) 0.00
B.3.4 plots
ASD .
Figure 15: Plot of the GIN $R^{2}$ results across post pooling layers probing for graph properties ($R^{2}$<0.1 have been hidden). (ABIDE dataset)
Figure 16: Plot of the GCN $R^{2}$ results across different layers probing for graph properties (ASD)
Figure 17: Plot of the GIN $R^{2}$ results across different layers probing for graph properties (ASD)
Figure 18: Plot of the GAT $R^{2}$ results across different layers probing for graph properties (ASD)
MDD .
Figure 19: Plot of the GIN $R^{2}$ results across different layers probing for graph properties ($R^{2}$<0.1 have been hidden). (REST-meta-MDD dataset).
Figure 20: Plot of the GCN $R^{2}$ results across different layers probing for graph properties (MDD)
Figure 21: Plot of the GIN $R^{2}$ results across different layers probing for graph properties (MDD)
Figure 22: Plot of the GAT $R^{2}$ results across different layers probing for graph properties (MDD)
B.3.5 Results ASD and MDD node properties
B.3.6 Node properties probing results
Table 25: Linear probing performance ( $R^{2}$ score on the test set) across models for various node properties (ASD dataset). Best Scores in Bold; Non-convergence indicated by —
GCN Layer degree closeness betweenness eigenvector clustering pagerank
x1 (GCN) 0.83 0.26 0.37 0.12
x2 (GCN) 0.73 0.29 0.02 0.37 0.16 0.43
x3 (GCN) 0.61 0.23 0.02 0.35 0.17 0.40
x4 (GCN) 0.53 0.19 0.03 0.31 0.17
out (GCN) 0.53 0.20 0.27 0.16
GAT Layer degree closeness betweenness eigenvector clustering pagerank
x1 (GAT) 0.55 0.07 0.05 0.32 0.28 0.17
x2 (GAT) 0.52 0.48 0.08 0.31 0.30 0.14
x3 (GAT) 0.47 0.55 0.29 0.29
x4 (GAT) 0.41 0.14 0.19 0.26
out (GAT) 0.35 0.50 0.12 0.21 0.23
GIN Layer degree closeness betweenness eigenvector clustering pagerank
x1 (GIN) 0.90 0.38 0.05 0.42 0.14 0.57
x2 (GIN) 0.89 0.24 0.12 0.40 0.16 0.59
x3 (GIN) 0.80 0.35 0.12 0.38 0.13 0.51
x4 (GIN) 0.82 0.42 0.17 0.36 0.11 0.70
out (GIN) 0.83 0.13 0.30 0.13 0.70
For ASD results, the strong presence of Page Rank is interesting. Regardless of this, without surprise it’s the degree that is consistently the highest node property as it prepare for global properties to aggregate.
Table 26: Linear probing performance ( $R^{2}$ score on the test set) across models for various node properties (MDD dataset). Best Scores in Bold; Non-convergence indicated by —
GCN Layer degree closeness betweenness eigenvector clustering pagerank
Layer 0 0.83 0.30 0.05 0.38 0.16 0.40
Layer 1 0.74 0.26 0.04 0.38 0.25
Layer 2 0.69 0.31 0.03 0.41 0.23
Layer 3 0.61 0.32 0.04 0.37 0.22
Layer 4 0.61 0.33 0.37 0.19
GAT Layer degree closeness betweenness eigenvector clustering pagerank
Layer 0 0.54 0.34 0.33 0.34 0.00
Layer 1 0.55 0.60
Layer 2 0.48 0.40 0.33 0.30 0.15
Layer 3 0.43 0.65 0.29 0.28
Layer 4 0.39 0.23 0.27
GIN Layer degree closeness betweenness eigenvector clustering pagerank
Layer 0 0.92 0.54 0.09 0.40 0.23 0.58
Layer 1 0.82 0.53 0.06 0.29 0.16 0.45
Layer 2 0.83 0.43 0.16 0.34 0.18 0.60
Layer 3 0.73 0.37 0.13 0.34 0.16 0.47
Layer 4 0.86 0.24 0.20 0.26 0.11 0.47
The MDD dataset shows similar results which are surely explained by the same arguments.
Appendix C Brain Imaging and GNNs
Our brain is a network, more precisely a complex network of functionally interconnected regions specialised in specific cognitive tasks, sharing information with each other. In the last three decades, the field of biological neuroscience and computational cognitive neuroscience have provided and incredible amount of knowledge on the role, function and biological structure of such regions of interests, aiming at better understanding both the biological organisation of the brain (which we can refer to as the ’hardware implementation’), the representation embedded in this hardware and the computational strategy employed to treat this kind of representation Marr (1984). In other terms, we got better at understanding how each region independently organises itself and processes and forms information (cite Connecting network science and information theory). The main problem for modern computational neuroscience consists of understanding the brain’s plasticity (how regions change over time), the inter-individual differences (how regions specialise differently between people) and how the brain integrates the information (how regions communicate with regard to each other).
For example we understand very well more basic brain structures like the cerebellum due to its high inter-individual similarity but we have a lot more difficulties modelling the prefrontal cortex which is so different from an individual to the other Kanai & Rees (2011); Gu & Kanai (2014); Mills et al. (2021). In other terms, we do understand well the brain operating in segregation but not so much in integration Aine (1995). Functional segregation refers to the distinct specialisation of anatomical brain regions and functional integration refers to the possible temporal dependencies between the activity of anatomically separated regions of the brain.
Because the representation of a system composed by agents and interactions among them by a complex network is an effective way to extract information on the nature and topology of such interactions, it makes a lot of sense to study the integration of the brain network through its temporal dependencies. Understanding the mathematical properties of such a network with regard to some functional state of the brain network therefore helps understanding how the integration system of the brain and its architecture are linked to ways of processing information. Using Marr’s paradigm to reformulate : understanding the functional communicative structure of the brain network helps understanding its algorithmic footprint. In terms of information theory, we could say that it helps understanding the relationship between topology and dynamics.
One way of accessing the brain activity is to use fMRI imaging. With fMRI measurements at ultra-high-field (3 Tesla, 7 Tesla or even 11 Tesla), hydrogen nuclei present in water and fat molecules align with the scanner’s powerful magnetic field. When radio waves briefly disturb this alignment, the nuclei return to their initial alignment with the magnetic field, this is known as the resonance and causes local changes in the magnetic field. These changes are detected by receiver coils. The collected data from these interactions enable the precise determination of the 3D locations of these events, in the so-called voxels, which can then be visualised. This process underlies the BOLD (Blood Oxygen Level Dependent) response, which is crucial for functional Magnetic Resonance Imaging (fMRI) as it reflects changes in blood flow and oxygenation associated with neuronal activity. We use the magnetic response of blood flow as a proxy for brain activity.
Then, relying on fMRI, we have several ways to study the functional connectivity of the brain. Functional connectivity is defined as the temporal dependence of neuronal activity patterns of anatomically separated brain regions Aertsen et al. (1989); Friston et al. (1993) and studies have shown that we could study functional connectivity between brain regions as the level of coactivation of functional MRI time-series Lowe et al. (1998; 2000). As a result, conceptualising the brain as an integrative network of functionally interacting brain regions offers a powerful framework for understanding large-scale neuronal communication. It provides a method for investigating how functional connectivity and information integration relates to human behaviour and how this organisation may be altered in neurodegenerative diseases Bullmore & Sporns (2009); Greicius et al. (2009).
To understand how a specific brain region interacts with others, researchers most often analyse its resting-state activity and use simple pearson correlation of time-series data of a region with the time-series data of all other brain regions, they create a functional connectivity map (fcMap), which visually represents the strength of these connections Biswal et al. (1997); Cordes et al. (2000). This is basically a matrix with value and we can understand it as a non relational data structure, in other terms, a graph.
More and more work in cognitive neurosciences explore the link between graph theory and connectomes (functional connectivity matrices) Farahani et al. (2019). By representing brain regions as nodes and their connections as edges, graph theory provides a powerful framework for analysing the structural and functional organisation of the brain. Notably, studies have begun to explore the link between structural properties of brain connectivity, as captured by connectomes, and the manifestation of neurological disorders such as Autism Spectrum Disorders (ASD) and Major Depressive Disorders (MDD). ASD, characterised by impairments in social communication and repetitive behaviours. MDD is characterised as a mood disorder marked by persistent sadness and loss of interest.
These findings highlight the potential of connectome analysis to elucidate the neurological underpinnings of NDs and pave the way for the development of novel diagnostic and therapeutic strategies. Studying the link between the brain’s Functional connectivity signature and behavioural quality of patients through probing learned embeddings of neural networks trained on classification tasks could thus be a promising avenue to help disentangle the gap between its segregational characteristics and the emergence Johnson (2002); Eccles (1994); Wang et al. (2015); Carroll & Parola (2024) of higher level behavioural quality.
However, if NDs result in alterations in brain functional and structural connections, as well as local and global connections Seeley et al. (2009); Wang et al. (2015); Pasquini et al. (2015); Stam et al. (2007), traditional deep learning models such as CNN and LSTM are difficult to fit to the connectivity of the brain Zhang et al. (2023). These long range dependencies, though, are well captured by the relational models defined previously in this thesis : Graph Neural Networks.
Definition : Psychiatric diagnosis can be regarded as a graph classification task. Given an input graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with node feature matrix $X$, GNNs employ the message-passing paradigm to propagate and aggregate the representations of information along edges to generate a node representation $h_{v}$ for each node $v\in\mathcal{V}$ and then explore the modelled human brains using graph methods to extract abnormal brain networks, subnetworks, and local connections Palop et al. (2006); Thomas et al. (2016).
Similarly to Zheng et al. (2023), a GNN can be formally defined through an aggregation function A and a combine function C such that $h_{v}^{(k)}$ is the node embedding of node $v$ at the $k$-th layer and $\mathcal{N}(v)$ is the set of neighbour nodes of $v$:
$\begin{gathered}a_{v}^{(k)}=\mathrm{A}^{(k)}\left(\left\{h_{u}^{(k-1)}:u\in\mathcal{N}(v)\right\}\right)\\ h_{v}^{(k)}=\mathrm{C}^{(k)}\left(h_{v}^{(k-1)},a_{v}^{(k)}\right)\end{gathered}$
In the context of connectomes, many studies have focused on the relationship between general intellectual ability and small-world characteristics in intrinsic functional networks for describing individual differences in general intelligence van den Heuvel & Hulshoff Pol (2010); van den Heuvel et al. (2009); Langer et al. (2012); Hilger et al. (2017). Better intellectual performance was associated with shorter characteristic path length, the nodal centrality of hub regions in the salience network, as well as the efficiency of functional integration between the frontal and parietal areas Jung & Haier (2007) In general, when connections between specialised brain regions are disrupted, even within localised areas, the result is often functional impairment. This impairment is linked to atypical integration of activity across distributed brain networks Ffytche & Catani (2005); Catani et al. (2005). Characterising this impairment through the use of GNN could be one application of our probing pipeline. So far, GNNs have achieved promising diagnostic accuracy on autism spectrum disorder (ASD) Rakhimberdina et al. (2020), schizophrenia Rakhimberdina & Murata (2020), bipolar disorder (BD) Yang et al. (2019) and MDD Zheng et al. (2023). We’ll focus on ASD and MDD. But here as in other graph related fields, research has highlighted we were lacking Interpretability Zheng et al. (2023).
For ASD, The contribution of rs-fMRI studies based on graph theory for autism exploration is important Redcay et al. (2013); Rudie et al. (2013); Di Martino et al. (2014); Keown et al. (2017); Kazeminejad & Sotero (2019). Studies have found increased short-range connections in ASD, particularly within sensory and association cortices. This local overconnectivity may contribute to the sensory sensitivities and restricted interests often seen in ASD. Conversely, long-range connections between distant brain regions tend to be reduced in ASD. This underconnectivity affects integration of information across brain networks. Based on this literature Farahani et al. (2019) we know that the modularity, clustering coefficient, and local efficiency are relatively reduced in ASD (i.e., inefficiency of information transmission in a particular module) while global communication efficiency is increased (shorter average path lengths). As another example, Redcay et al. (2013) observed an increase in betweenness centrality and local connections by analysing the prefrontal brain areas in adolescents with ASD.
In the node property level, we would expect betweenness centrality to be one of the major properties linked with ASD. In the graph level level, we would thus expect the clustering coefficient, the degree to which connected nodes in the brain network are clustered together indicating increased local processing and functional segregation and over-connectivity in local brain regions. We would also expect the characteristic path length to be disrupted, the average shortest path length between all pairs of nodes in the network, suggesting differences in global information transfer efficiency. And small-worldness (SW) which quantifies the balance between local clustering and global integration. Atypical SW in ASD may reflect disrupted optimal network organisation imbalance between local and global processing. We would expect these properties to be critical in our GNNs embeddings trained on classification tasks.
For patients with MDD, several studies have reported topological changes in human brain connectome, including a loss of the small-world network Ye et al. (2015); Achard & Bullmore (2007)] and a significant reorganisation of the community structure Zhang et al. (2011); Leistedt et al. (2009); Lord et al. (2012). In general, MDD patients exhibit increased global and local clustering coefficients, indicating a higher degree of local interconnectedness and efficiency in information processing. Moreover, increased modularity in MDD patients indicated that there were relatively less inter-modular edges and more intra-modular edges, which may also be associated with the disruptions in emotion regulation by decreasing communications between the Default Mode Network (DMN) and the Cognitive Control Network (CCN) Ye et al. (2015). We would thus expect that classifying FC matrices with regard to MDD should use more clustering coefficient, clusterization properties and modularity measures than random (like the presence of motifs like the number small clusters, squares or triangles).